Find the amount that results from the investment.
$480 invested at 16% compounded quarterly after a period of 4 years.
Possible answer:
a. $864.45
b. $419.03
c. $869.11
d. $899.03

Answer by rfer(16322)

FV=480(1+0.16/4)^4*4
FV=480(1.04)^16
FV=480*1.87298
FV=$899.03 (D)